In 2017, the entire fleet of light‑duty vehicles sold in the United States by each manufacturer must emit an average of no more than 84 milligrams per mile (mg/mi) of nitrogen oxides (NOX) and nonmethane organic gas (NMOG) over the useful life ( 150,000 miles of driving) of the vehicle. NOX + NMOG emissions over the useful life for one car model vary Normally with mean 78 mg/mi and standard deviation 6 mg/mi.

(a) What is the probability that a single car of this model emits more than 84 mg/mi of NOX + NMOG? (Enter your answer rounded to four decimal places.)

(b) A company has 3636 cars of this model in its fleet. What is the probability that the average NOX ++ NMOG level ?¯x¯ of these cars is above 8484 mg/mi? (Enter your answer rounded to four decimal places.)
Answer #1

Solution :

Given that ,

mean = \mu = 78

standard deviation = \sigma = 6

P(x >84 ) = 1 - P(x < 84)

= 1 - P[(x - \mu ) / \sigma < (84 - 78) / 6]

= 1 - P(z <1 )

Using z table,

= 1 -0.8413

=0.1587

(B)n = 36

\mu\bar x = 78

\sigma\bar x = \sigma / \sqrt n =6 / \sqrt 36= 1

P(\bar x >84 ) = 1 - P(\bar x <84 )

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (84 - 78) / 1]

= 1 - P(z <6 )

Using z table,    

= 1 - 1

= 0